OpenStudy (jkbo):

Please help Which determinant has a value of 2? a.det [-7 9] [1 1] b.det [7 9] [1 1] c.det [7 -9] [1 1] d.det [-7 -9] [1 1]

OpenStudy (anonymous):

\[-7\times 1-1\times 9=-16\] so not the first one

OpenStudy (jkbo):

Its D.

OpenStudy (jkbo):

thank you very

OpenStudy (anonymous):

yw